# Tropicalization through the lens of Łukasiewicz logic, with a topos theoretic perspective

### v1

@@ 3.1 $\theta$ invariant of $V(C)$-algebras
Since $\theta(A)$ is the zerosets of an MV-polynomial, that is an algebraic variety of the $V(C)$, it is invariant with respect to isomorphic MV-algebras. Note that for every $A\in V(C)$,
  • $B(A)\subseteq\theta(A)$ (in particular, $0,1\in\theta(A)$),
  • $Rad(A)\subseteq\theta(A)$
  • $co-Rad(A)\cap\theta(A)=\{1\}$.
Lemma 3.1.
Let $A$ be an MV-algebra in $V(C)$. Let $x\in A$. The following are equivalent.
  • $x\geq 2x^{2}$, that is, $x\in\theta(A)$.
  • $x$ is a sum of a boolean and an infinitesimal.
  • for every prime ideal $P$, $x\ mod\ P$ is infinitesimal or $1$.
@@
We first prove that the second condition implies the first. Since every algebra in $V(C)$ is subdirects product of linearly ordered $V(C)$-algebras, it is enough to consider linearly ordered $V(C)$-algebras. In this case, the boolean part of $x$ is 00 or $1$, so $x$ is either an infinitesimal or $1$. In both cases, we have $x\geq 2x^{2}$.
Conversely, we know that every element $x$ of an MV-algebra in $V(C)$ decomposes as $x=(b\oplus\epsilon)\ominus\tau$, where $b$ is boolean and $\epsilon,\tau$ infinitesimals such that $\epsilon\wedge\tau=0$. Since every algebra in $V(C)$ is subdirects product of linearly ordered $V(C)$-algebras $L_{i}$ for $i$ in an index set $I$, we can consider the $i$-th component of the expression: $x_{i}=(b_{i}\oplus\epsilon_{i})\ominus\tau_{i}$. If $b_{i}=0$ then we must have $\tau_{i}=0$. If $b_{i}=1$, then from the hypothesis we infer $x_{i}\geq 2x_{i}^{2}$ and this implies $\tau_{i}=0$. Summing up, $\tau=0$ and $x=b\oplus\epsilon$.
The first condition implies the third since it is essentially equational. Conversely, if the third condition holds, then $x\geq 2x^{2}$ holds modulo any prime $P$ of $A$, so it holds in $A$.
@@
In general $\theta(A)$ is not closed under negation. In fact we have:
Lemma 3.3.
Let $A\in V(C)$. The following are equivalent.
  • $\theta(A)$ is closed under negation.
  • $A=\theta(A)$.
  • $\theta(A)$ is an MV-algebra.
  • $A$ is a Boolean algebra.
Proof.
If $A=\theta(A)$ then infinitesimals cannot exist in $A$, otherwise coinfinitesimals would exist in $\theta(A)$, but coinfinitesimals never satisfy $x\geq 2x^{2}$ unless $x=1$.
If $A$ is a Boolean algebra then $x=x^{2}=2x^{2}$ for every $x\in A$, hence $A=\theta(A)$. ∎
4 Semirings, $\theta$ reducts and $\ell$-Bisemirings
In this section we will analyze the algebraic aspects of $\theta$ invariant. We see that the algebraic variety $\theta(A)$ can be endowed with a rich algebraic structure. This opens the door to classifications of these varieties.
@@ 4 Semirings, $\theta$ reducts and $\ell$-Bisemirings
    $(S,\wedge,0,1,\oplus)$ is an idempotent commutative semiring; $(S,\vee,0,1,\odot)$ idempotent commutative semiring.
@@
Definition 4.2.
A Coupled $\ell$-Bisemirings is a two sorted first order structure $(S,T,\nu)$ where
  1. 1.
    $S$ and $T$ are $\ell$-Bisemirings;
  2. 2.
    $\nu$ is an isomorphism between $S$ and $T$ swapping $\vee$,$\wedge$ , ∧ and $\oplus$,$\odot$ , ⊙.
We denote by $C\ell BS$ the category of Coupled $\ell$-Bisemirings and their morphisms (as two sorted structures of first order logic). Namely, a morphism $(S,T,\nu)\to(S^{\prime},T^{\prime},\nu^{\prime})$ is a pair of $\ell$-Bisemiring homomorphisms $(f,g)$ such that $f:S\to S^{\prime}$, $g:T\to T^{\prime}$, and $\nu^{\prime}\circ f=g\circ\nu$.
In the previous section we introduced functors $\theta$ and $\theta^{*}$ from $V(C)$-algebras to $Set$. We observe that $\theta(A)$ and $\theta^{*}(A)$ are algebraic varieties in the sense of algebraic geometry, despite the structure used in defining them is the one of MV-algebras instead of rings or fields. The negation in $A$ is an anti isomorphism between $\theta(A)$ and $\theta^{*}(A)$.
Let us define $\Theta(A)=(\theta(A),\theta^{*}(A),\neg)$.
@@ A Coupled $\ell$-Bisemirings is a two sorted first order structure $(S,T,\nu)$ where
For every $A\in V(C)$, the structure $\Theta(A)$ belongs to $C\ell BS$. The map sending $A$ to $\Theta(A)$ is functor from $V(C)$ to $C\ell BS$.
@@
The same holds for $\theta^{*}$.
In particular, $(\theta^{*}(A),\wedge,\oplus,1,0)$ is again a commutative, additively idempotent semiring.
On the other hand, with every MV-algebra $A$ we can associate a Coupled $\ell$-Bisemiring by considering $(A,A,id_{A})$. Thus we have a functor from the category of MV-algebras to $C\ell BS$ and a different functor from the category of $V(C)$ to $C\ell BS$.
The structure $(\theta(A),\oplus,\odot,\vee,\wedge,0,1)$ is a Coupled $\ell$-Bisemiring: we want now to characterize such structures.
@@
Let $(S,T,\nu)$ be a Coupled $\ell$-Bisemiring. Then $S=\theta(A)$ for some $A\in V(C)$ if and only if:
  • $(Inf(S),\oplus,\odot,\wedge,\vee)$ is the radical of an MV-algebra, where $Inf(S)$ is the set of $x\in S$ such that $x^{2}=0$.
  • $(BOOL(S),\oplus,\odot)$ is a Boolean algebra, where $\vee$=$\oplus$ = ⊕ and $\wedge$=$\odot$ = ⊙ and $Bool(S)$ is the set of $x$ such that $x^{2}=x$.
Proof.
$\theta(A)$ has the required properties for every $A\in V(C)$.
Conversely, to obtain $A$ from $S$ it is enough take an MV-algebra $A\in V(C)$ such that $B(A)=Bool(S)$ and $Rad(A)=Inf(S)$. One such MV-algebra is built in the previous lemma. ∎
By specializing the result to perfect MV-algebras we have:
Corollary 4.1.
Let $(S,T,\nu)$ be a Coupled $\ell$-Bisemiring. Then $S=\theta(A)$ for some perfect MV-algebra $A$ if and only if:
  • $(Inf(S),\oplus,\odot,\wedge,\vee)$ is the radical of an MV-algebra, where $Inf(S)$ is the set of $x\in S$ such that $x^{2}=0$.
  • $(Bool(S),\oplus,\odot)$ is a two element Boolean algebra, where $\vee$=$\oplus$ = ⊕ and $\wedge$=$\odot$ = ⊙ and $Bool(S)$ is the set of $x$ such that $x^{2}=x$.
It would be interesting to classify all MV-algebras with a given $\theta$. Probably one should consider actions of boolean algebras on monoids (a link with the arithmetic site of [12]?).
We have seen that the algebraic variety $\theta(A)$, with $A\in V(C)$, have a lot of algebraic structure. This is confirmed by the following operations, all available in $\theta(A)$:
  • $bool:A\to B(A)$, $bool(x)=2x^{2}$ (the boolean part of $x$)
  • $inf:A\to Rad(A)$, $inf(x)=x\wedge\neg x$ (the infinitesimal part of $x$)
  • $neg_{b}:A\to B(A)$, $\neg_{b}(x)=\neg(bool(x))$.
5 Tropicalization of Perfect MV-algebras
The aim of this section is to show how a generalization of the concept of tropicalization of algebraic structures makes it possible allows one to discover interesting inter-relationship between categories of algebraic models of $\L$ukasiewicz logics and their algebraic geometries through a tropicalization functor.
@@ 5 Tropicalization of Perfect MV-algebras
We explore the role of $\theta$ functor in composition with other functors acting in tropical geometry, and more, its role in moving from $V(C)$-algebraic varieties into tropical algebraic varieties.
@@
Let $TOP$ be the category of positive cones of $\ell$-groups with a top, where morphisms preserve the positive cone and the top. Let $BOT$ be the category of negative cones of $\ell$-groups with a bottom, where morphisms preserve the negative cone and the bottom. Let $PERF$ be the category of perfect MV-algebras.
@@
$\theta$ extends to an equivalence between $PERF$ and $TOP$. $\theta^{*}$ extends to an equivalence between $PERF$ and $BOT$.
Proof.
Let $P$ be a perfect MV-algebra. Then $\theta(P)=Rad(P)\cup\{1\}$ and $Rad(P)$ is the positive cone of the $\ell$-group $\Delta^{-1}(P)$. Then it is easy to check that the claim of the theorem holds. ∎
Let $G$ be an abelian $\ell$-group. The $\Delta$ functor maps $G$ in a perfect MV-algebra $\Delta(G)$. Let $\ell\cal G$ be the category of abelian $\ell$-groups. Applying the functors $\theta,\theta^{*}$ we obtain:
Theorem 5.2.
@@
$\theta\circ\Delta$ is a categorial equivalence from $\ell\cal G$ to $TOP$. $\theta^{*}\circ\Delta$ is a categorial equivalence from $\ell\cal G$ to $BOT$.
Now if $G$ is an abelian l-group, we define its tropicalization $Trop(G)$ as the additively idempotent semifield obtained by adding a point $\infty$- ∞ to $G$. The tropicalization can be considered as a functor from $\ell\cal G$ to the category of idempotent semirings. The inverse functor takes an additively idempotent semifield $S$ and maps it to the abelian l-group $Detrop(S)$ obtained by deleting the zero of $S$.
So, given an additively idempotent semifield $S$, we can define two functors $\mathbb{F},\mathbb{F^{*}}$ by a triple composition of functors
$\mathbb{F}(S)=\theta(\Delta(Detrop(S))),$
$\mathbb{F^{*}}(S)=\theta^{*}(\Delta(Detrop(S))),$
Theorem 5.3.
$\mathbb{F}$ is an equivalence from the category of additively idempotent semifields to $TOP$. $\mathbb{F^{*}}$ is an equivalence from the former category to $BOT$.
So the functor $\mathbb{F}$ is an equivalence, and gives a direct link between tropical structures ( additively idempotent semifields) and our $\theta$ structures over Perfect MV-algebras. The construction is functorial, and is more general than dequantization, see [38], in the sense that dequantization leads only to the construction of single structures like the (max,plus)-semiring.
So, the above theorem seems to be the answer to what we propose in the introduction, where, namely, we proposed to build up a functorial tropicalization of an algebraic structure. Indeed, we see the elements of $TOP$ as tropicalizations of perfect MV-algebras. It remains to tropicalize in a reasonable sense more general MV-algebras.
6 MV-algebras and the topos $\widehat{\mathbb{N}^{\times}}$
In this section, we will show how a logic admitting only truth values infinitesimally close to boolean values, can surprisingly exhibit models that are functorially connected to a category of points of a non-commutative geometry.
We are inspired by [12] where the authors apply topos theory and tropical geometry for a possible road to Riemann Hypothesis. The hope is that, as schemes opened the way for Fermat’s theorem, topoi open the way for Riemann hypothesis. In [12] we found the crucial construction presented here by
Definition 6.1.
The points of the topos $\widehat{\mathbb{N}^{\times}}$ form a category which is studied in [12]. We aim to establish some model theoretic properties of classes of perfect MV-algebras. Then, we will specialize the obtained results to the category of $Perf^{\mathbb{Q}}$, below defined, and proceeding along this road to get model theoretic properties of the class of points of $\widehat{\mathbb{N}^{\times}}$.
Indeed we aim to give a MV-algebraic characterization of the points of the presheaf topos $\widehat{\mathbb{N}^{\times}}$ and the logical consequences of this characterization.
We associate points of $\widehat{\mathbb{N}^{\times}}$ to theories in $L_{p}$, the extension of Lukasiewicz logic with the equation defining the variety $V(C)$.
Definition 6.2.
Call an ordered monoid subtractive if $x\leq y$ holds if and only if there is $z$ with $y=x+z$.
Let us call $Canc^{\mathbb{Q}}$ the category of cancellative subtractive lattice ordered abelian submonoids of $\mathbb{Q}^{+}$ with an extra bottom element.
The paper [8] gives an equivalence between abelian $\ell$-groups and cancellative subtractive lattice ordered abelian monoids with bottom element. The idea is to associate to an $\ell$-group its positive cone plus an infinite point. By restricting the equivalence to the subgroups of $\mathbb{Q}$, we obtain an equivalence between $Ab^{\mathbb{Q}}$ and $Canc^{\mathbb{Q}}$. Summing up we have:
Theorem 6.1.
The categories $Ab^{\mathbb{Q}}$, $TOP^{\mathbb{Q}}$, $Canc^{\mathbb{Q}}$ and $Perf^{\mathbb{Q}}$ are pairwise equivalent.
Now we give some translation lemmas. Note that if $A$ is an MV-algebra then $Rad(A)$ is closed with respect to operations $\oplus$ and $\ominus$.
Lemma 6.1.
Let $A$ be a perfect MV-algebra and let $\phi$ be a formula of the language of MV-algebras. Then $A\models\phi(x_{1},\ldots,x_{n})$ if and only if $Rad(A)\models Rad(\phi)(y_{1},\ldots,y_{n})$, where $Rad(\phi)$ is a formula in the language of $\oplus$,$\ominus$ , ⊖, and $y_{i}=x_{i}$ if $x_{i}$ has type Rad, and $y_{i}=\neg x_{i}$ otherwise.
Proof.
Since $A$ is perfect, $A$ is the union of the radical and the coradical. So we can suppose that the variables of $\phi$ have type $Rad$ or $Corad$.
The type of a term (rad or corad) is defined inductively as follows. 00 has type rad. $\neg t$ has type rad if and only if $t$ has type corad. Finally $t\oplus u$ has type rad if both $t,u$ have type rad, and corad otherwise.
The radical is equipped with the functions $\oplus$,$\ominus$ , ⊖.
We prove inductively that every term of type Rad is a combination of $\oplus$,$\ominus$ , ⊖ of variables of type rad and negated variables of type corad.
Suppose a term $t$ is of type Rad.
If $t$ is a variable then $t$ has the desired form.
If $t=\neg u$ and $u$ is a variable then $t$ has the desired form.
If $t=u\oplus v$ then $u,v$ have type rad and we are done.
If $t=\neg\neg u$ then $u$ has the desired form.
Suppose $t=\neg(u\oplus v)=\neg u\odot\neg v$. If $u,v$ have type rad then $t$ has type corad, which is a contradiction.
If $u,v$ have type corad then $t=0$.
If $u$ has type rad and $v$ has type corad then $\neg v$ has type rad and $t=(\neg v)\ominus u$ which has the desired form. Similarly we proceed if $u$ is corad and $v$ is rad.
We can suppose that every term in $\phi$ has type rad. In fact, suppose $t=u$ have different types. Then $t=u$ can be replaced by false. If $t=u$ occurs in $\phi$ and $t,u$ have type corad, then $t=u$ can be replaced by $\neg t=\neg u$. So the only case remaining is that only terms of type rad occur in $\phi$.
Now the formula $\exists x:corad.\phi(\neg x)$ can be replaced by $\exists x:rad.\phi(x)$. In this way $\phi$ is transformed in a formula whose variables are all of type rad, and all terms are combinations of $\oplus$ and $\ominus$ of variables.
Going from radicals to groups we get:
Lemma 6.2.
Take a formula $\phi$ in the language of $\oplus$,$\ominus$ , ⊖. There is a formula $\Delta_{0}(\phi)$ in the language of ordered groups, effectively built from $\phi$, such that for every lattice ordered abelian group $G$, we have $Rad(\Delta(G))\models\phi(x_{1},\ldots,x_{n})$ if and only if $G\models\Delta_{0}(\phi)(x_{1},\ldots,x_{n})$.
Proof.
This is because in $G$ we have $x\oplus y=x+y$, $x\ominus y=0\vee(x-y)$. ∎
Now by combining the translations $\Delta_{0}$ and $Rad$ we obtain:
Theorem 6.2.
Take a formula $\phi$ in the language of MV-algebras whose variables range over the radical or the coradical. There is a formula $\Delta(\phi)$ in the language of ordered groups, effectively built from $\phi$, such that for every lattice ordered abelian group $G$, we have $\Delta(G)\models\phi(x_{1},\ldots,x_{n})$ if and only if $G\models\Delta(\phi)(y_{1},\ldots,y_{n})$, where $y_{i}=x_{i}$ if $x_{i}$ ranges over the radical, and $y_{i}=\neg x_{i}$ otherwise.
@@ 6 MV-algebras and the topos $\widehat{\mathbb{N}^{\times}}$
We will see that the duality $\Delta$ can be used to transfer logical information between abelian $\ell$-groups and perfect MV-algebras, in particular, regarding first order logic. Something similar happens for the duality $\Gamma$, but not quite for first order logic, rather geometric logic in the sense of topos theory, since both MV-algebras and $\ell$ groups with strong units are axiomatized in geometric logic.
Here we recall that a class of models is geometric if it is axiomatized by geometric sequents in the sense of topos theory, see [6].
Geometric logic is a fragment of infinitary first order logic of a constructive flavor where one allows infinite disjunctions, finite conjunctions and existential quantifiers (no negation).
We start by investigating the first order theory of the non trivial subgroups of the rationals as ordered groups. Then via the $\Delta^{\mathbb{Q}}$ functor we can transfer the obtained results to the first order theory of $Perf^{\mathbb{Q}}$ and to points of such a natural topos $\widehat{\mathbb{N}^{\times}}$.

### v2

@@ 3 Rump tropicalization and its generalization via $V(C)$ algebras: the functor $\theta$
For example, for $C=\Gamma(\mathbb{Z}\stackrel{{\scriptstyle\mbox{\tiny{\emph{lex}}}}}{{\times}}\mathbb{Z},(1,0)$ we have $\theta(C)=\{(0,n)\mid n\in\mathbb{N}\}\cup\{(1,0)\}$. Since $\theta(A)$ is the zeroset of an MV-polynomial, that is an algebraic variety of the $V(C)$, it is invariant with respect to isomorphic MV-algebras. In the following we collect some properties of $\theta$ and $\theta^{*}$:
Proposition 3.1.
For every $A\in V(C)$, $x\in\theta^{*}(A)$ if and only if $\neg x\in\theta(A)$. Hence $\theta(A)$ is closed under negation if and only if $\theta(A)=\theta^{*}(A)$.
@@
$x\in\theta^{*}(A)$ iff $x\leq 2x^{2}$, and applying the negation, $\neg x\geq\neg(2x^{2})$ iff $\neg x\geq 2(\neg x)^{2})$ iff $\neg x\in\theta(A)$.
Proposition 3.2.
For every $A\in V(C)$, $B(A)\subseteq\theta(A)$ (in particular, $0,1\in\theta(A)$) and $Rad(A)\subseteq\theta(A)$. Further, $co-Rad(A)\cap\theta(A)=\{1\}$ and, for every $x\in A$, $x\in\theta(A)$ if and only if $x$ is a sum of a boolean and an infinitesimal.
@@
In general $\theta(A)$ is not closed under negation. Indeed, suppose $\theta(A)$ is closed under negation. Then infinitesimals cannot exist in $A$, otherwise infinitesimals would exist in $\theta(A)$ and, by negation, coinfinitesimals would exist in $\theta(A)$. But coinfinitesimals never satisfy $x\geq 2x^{2}$ unless $x=1$. So $A$ is a semisimple MV-algebra in $V(C)$, hence $A$ is a Boolean algebra and $A=\theta(A)$, in particular $\theta(A)$ is an MV algebra.
@@
    $(S,\wedge,0,1,\oplus)$ is an idempotent commutative semiring;
  3. 3.
    $(S,\vee,0,1,\odot)$ is an idempotent commutative semiring.
@@
We denote by $\ell BS$ the category of $\ell$-Bisemirings and their morphisms.
@@
For every $A\in V(C)$, the structure $\theta(A)$ belongs to $\ell BS$ and the map sending $A$ to $\theta(A)$ is functor from $V(C)$ to $\ell BS$. The same holds for $\theta^{*}$.
@@
A similar proof holds for $\theta^{*}$.
In a sense, negation gives a duality between the functors $\theta$ and $\theta^{*}$.
@@
Let $S$ be an $\ell$-Bisemiring. Then $S=\theta(A)$ for some $A\in V(C)$ if and only if:
  • $(Inf(S),\oplus,\odot,\wedge,\vee)$ is the radical of an MV-algebra, where $Inf(S)$ is the set of $x\in S$ such that $x^{2}=0$ (infinitesimal elements).
  • $(BOOL(S),\oplus,\odot)$ is a Boolean algebra, where $\vee$=$\oplus$ = ⊕ and $\wedge$=$\odot$ = ⊙ and $Bool(S)$ is the set of $x$ such that $x^{2}=x$ (Boolean elements).
In particular, if $S=\theta(A)$ for a perfect MV-algebra $A$, then $BOOL(S)=\{0,1\}$.
3.1 Tropicalization of Perfect MV-algebras
@@
We explore the role of $\theta$ in composition with other functors acting in tropical geometry, and more, its role in moving from $V(C)$-algebraic varieties into tropical algebraic varieties.
@@
Let $TOP$ be the category of positive cones of $\ell$-groups with a top, where morphisms preserve the positive cone and the top, $PERF$ be the category of perfect MV-algebras and $\ell\cal G$ be the category of abelian $\ell$-groups.
For any perfect MV-algebra $P$, $\theta(P)=Rad(P)\cup\{1\}$ and $Rad(P)$ is the positive cone of the $\ell$-group $\Delta^{-1}(P)$. On the other hand, the $\Delta$ functor maps every abelian $\ell$-group $G$ in a perfect MV-algebra $\Delta(G)$. We can extend such correspondences to equivalences of categories:
@@
$\theta$ is an equivalence functor between $PERF$ and $TOP$. $\theta\circ\Delta$ is a categorial equivalence from $\ell\cal G$ to $TOP$.
@@
If $G$ is an abelian $\ell$-group, we define its tropicalization $Trop(G)$ as the additively idempotent semifield obtained by adding a point $\infty$- ∞ to $G$. The tropicalization can be considered as a functor from $\ell\cal G$ to the category of idempotent semirings. The inverse functor takes an additively idempotent semifield $S$ and maps it to the abelian $\ell$-group $Detrop(S)$ obtained by deleting the zero of $S$.
So, given an additively idempotent semifield $S$, we can define the functor $\mathbf{F}$ by a triple composition of functors
$\mathbf{F}(S)=\theta(\Delta(Detrop(S))),$
Theorem 3.3.
$\mathbf{F}$ is an equivalence from the category of additively idempotent semifields to $TOP$.
So the functor $\mathbf{F}$ is an equivalence, and gives a direct link between tropical structures ( additively idempotent semifields) and our $\theta$ structures over Perfect MV-algebras. The construction is functorial.
So, the above theorem build up a functorial tropicalization of algebraic structures related with Łukasiewicz logic. Indeed, we see the elements of $TOP$ as tropicalizations of perfect MV-algebras. It remains to tropicalize in a reasonable sense more general MV-algebras.
4 MV-algebras and the topos $\widehat{\mathbb{N}^{\times}}$
We are inspired by [8] where the authors apply topos theory and tropical geometry for a possible road to Riemann Hypothesis. The hope is that, as schemes opened the way for Fermat’s theorem, topoi open the way for Riemann hypothesis.
We recall the reader that a topos is a mathematical structure that generalizes the concept of a space, extending the ideas of set theory and logic into a categorical framework. Originating from Grothendieck’s work in algebraic geometry, topoi have found applications in diverse fields such as logic, computer science, and topology. At its core, a topos can be viewed as a category that behaves like the category of sets, satisfying specific axioms that make it a generalization of a ‘‘universe of discourse."
One of the simplest examples of a topos is a presheaf category, denoted as $\text{Set}^{\mathcal{C}^{\text{op}}}$, where $\mathcal{C}$ is a small category. Objects in this category are functors from $\mathcal{C}^{\text{op}}$ to the category of sets, and morphisms are natural transformations between these functors. A presheaf hence assigns sets to objects in $\mathcal{C}$ and specifies how these sets transform along morphisms of $\mathcal{C}$.
A key feature of topoi is the concept of points, which abstractly generalize the idea of geometric points in a space. A point of a topos $\mathcal{E}$ is defined as a geometric morphism $\mathsf{Set}\to\mathcal{E}$, where $\mathsf{Set}$ is the category of sets. Such a morphism consists of an adjoint pair of functors that preserve the essential structure of the topos. Points play a significant role in connecting the categorical structure of a topos to more classical notions of geometry and logic.
In summary, topoi unify geometric and logical perspectives, presheaf categories provide an intuitive entry point into their study, and points of a topos link them to tangible, classical ideas. Together, they form a rich and versatile framework for exploring abstract and applied mathematics.
In [8] we found the crucial construction from which we start our investigation:
Definition 4.1.
The points of the topos $\widehat{\mathbb{N}^{\times}}$ form a category which is studied in [8].
We aim to give a MV-algebraic characterization of the points of the presheaf topos $\widehat{\mathbb{N}^{\times}}$ and to investigate the logical consequences of this characterization.
We start by investigating the first order theory of the non trivial subgroups of the rationals as ordered groups. Then via the $\Delta^{\mathbb{Q}}$ functor we can transfer the obtained results to the first order theory of $Perf^{\mathbb{Q}}$ and to points of the topos $\widehat{\mathbb{N}^{\times}}$.